# Calibrating dimension reduction hyperparameters in the presence of noise

### v3

@@ Abstract
The goal of dimension reduction tools is to construct a low-dimensional representation of high-dimensional data. These tools are employed for a variety of reasons such as noise reduction, visualization, and to lower computational costs. However, there is a fundamental issue that is highly discussed in other modeling problems, but almost entirely ignored in the dimension reduction literature: overfitting. If we interpret data as a combination of signal and noise, prior works judge dimension reduction techniques on their ability to capture the entirety of the data, i.e. both the signal and the noise. In the context of other modeling problems, techniques such as feature-selection, cross-validation, and regularization are employed to combat overfitting, but no such precautions are taken when performing dimension reduction. In this paper, we present a framework that models dimension reduction problems in the presence of noise and use this framework to explore the role perplexity and number of neighbors play in overfitting data when applying t-SNE and UMAP. More specifically, we show previously recommended values for perplexity and number of neighbors are too small and tend to overfit the noise. We also present a workflow others may use to calibrate hyperparameters in the presence of noise.
@@ 1 Introduction
In recent years, non-linear dimension reduction techniques have been growing in popularity due to their usefulness when analyzing high-dimensional data. To date, they are the most effective tools for visualizing high-dimensional data. Visualization plays a key role in identifying and understanding relationships within the data. It exposes patterns, trends, and outliers that can’t be discerned from the raw data. The most popular non-linear dimension reduction technique is t-distributed stochastic neighbor embedding (t-SNE, [17]).
@@
Since the introduction of t-SNE, hyperparameter calibration has proven to be a difficult task. The most crucial hyperparameter, known as perplexity, controls how large a neighborhood to consider around a point when determining its location. Perplexity calibration is so troublesome, that perplexity-free versions of t-SNE have been proposed [6]. It is also an extremely important task, since t-SNE is known to produce unfaithful results when mishandled [9]. The original authors suggested values between 5 and 50 [17], while recent works have suggested perplexities as large as one percent of the sample size [10]. [3] studied the inverse relationship between perplexity and Kullback-Leibler divergence to design an automatic calibration process that “generally agrees with experts’ consensus.” Manual tuning of perplexity requires a deep understanding of the t-SNE algorithm and general knowledge of the data’s structure. When the data’s structure is available, we can visualize the results and choose the perplexity that best captures the hypothesized structure. In supervised problems, for example, we look for low-dimensional representations that cluster according to the class labels. For unsupervised problems, however, the structure is often unknown, so we cannot visually assess each representation. In these cases, we must resort to quantitative measures of performance to understand how the well the low-dimensional representation represents the high-dimensional data. While this strategy is heavily discussed in the machine learning literature, prior works disregard the possibility of overfitting when measuring performance.
@@
In this paper, we present a framework for studying dimension reduction methods in the presence of noise (Section 3). We then use this framework to calibrate t-SNE and UMAP hyperparameters in both simulated and practical examples to illustrate how the disregard of noise leads to miscalibration (Section 4). We also discuss how other researchers may use this framework in their own work (Section 5).
@@ 2 Background
Thus for each point $x_{i}$, we have a probability distribution $P_{i}$ that quantifies the similarity of every other point to $x_{i}$. The scale of the Gaussian kernel, $\sigma_{i}$, is chosen so that the perplexity of the probability distribution $P_{i}$, in the information theory sense, is equal to a pre-specified value also named perplexity,
@@
Intuitively, perplexity controls how large a neighborhood to consider around each point when approximating the topological structure of the data [17]. As such, it implicitly balances attention to local and global aspects of the data with high values of perplexity placing more emphasis on global aspects. For the sake of computational convenience, t-SNE assumes the directional similarities are symmetric by defining
@@
The main downsides of t-SNE are its inherit randomness and sensitivity to hyperparameter calibration. The minimization of the KL divergence is done using gradient descent methods with incorporated randomness to avoid stagnating at local minima. As a result, the output differs between runs of the algorithm. Hence, the traditional t-SNE workflow often includes running the algorithm multiple times at various perplexities before choosing the best representation.
@@ 2.2 UMAP
Uniform Manifold Approximation and Projection (UMAP, [12]) is another nonlinear dimension reduction method that has been rising in popularity. Like t-SNE, UMAP is a powerful tool for visualizing high-dimensional data that requires user calibration. The architecture of the UMAP algorithm is similar to that of t-SNE’s — high-dimensional similarities are computed and the resulting representation is the set of low-dimensional points whose low-dimensional similarities best match the high-dimensional similarities. The cost function and the formulas for high/low-dimensional similarities differ from t-SNE. See [12] for details.
UMAP has a hyperparameter called n_neighbors that is analogous to t-SNE’s perplexity. It determines how large a neighborhood to use around each point when calculating high-dimensional similarities. The original authors make no recommendation for optimal values of n_neighbors, but their implementation defaults to n_neighbors = 15 [12].
@@ 3.1 Dimension Reduction Framework
Prior works quantitatively measure how well low-dimensional representations match the high-dimensional data. However, if we view the original data as a composition of signal and noise, we must not reward capturing the noise. This would lead to overfitting. Therefore, we should be comparing the low-dimensional representation against the signal underlying our data, rather than the entirety of the data.
Suppose the underlying signal of our data is captured by an $r$-dimensional matrix $Y\in\mathbb{R}^{n\times r}$. In the context of dimension reduction, the underlying signal is often lower dimension than the original data. Let $p\geq r$ be the dimension of the original data set, and let $\textrm{Emb}:\mathbb{R}^{r}\to\mathbb{R}^{p}$ be the function that embeds the signal in data space. Define $Z=\textrm{Emb}(Y)$ to be the signal embedded in data space. We then assume the presence of random error. The original data can then be modeled by
$Z+\epsilon\textrm{ for }\epsilon\sim N_{p}(0,\Sigma).$
The dimension reduction method $\varphi$ is applied to $Z+\epsilon$ to get a low-dimensional representation $X\in\mathbb{R}^{n\times q}$. See Figure 1.
{tikzcd}
@@ 3.2 Reconstruction Error Functions
The remaining piece is a procedure for measuring dimension reduction performance. Suppose we have a reconstruction error function $f(D_{1},D_{2})$ that quantifies how well the data set $D_{2}$ represents the data set $D_{1}$. Prior works like [4], [7], [9], and [10] use various reconstruction error functions to quantify performance; only, they study $f(Z+\epsilon,X)$ to measure how well the constructed representation $X$ represents the original data $Z+\epsilon$. We argue it is more appropriate to compare $X$ against the underlying signal $Y$ by examining $f(Y,X)$.
Prior works in dimension reduction have suggested various quantitative metrics for measuring dimension reduction performance. In line with recent discussions of perplexity ([10] and [4]), we focus on two different metrics — one that measures local performance and one that measures global performance.
@@
For global performance, we use Shepard goodness [7]. Shepard goodness is the Kendall correlation (a rank-based correlation) between high and low-dimensional inter-point distances,
$f_{\textrm{Shep}}(D_{1},D_{2})=\sigma_{\textrm{Kendall}}(||z_{i}-z_{j}||^{2},||\varphi(z_{i})-\varphi(z_{j})||^{2}).$
@@ 3.3 Using this framework
When using our framework to model examples, three components must be specified: $Z+\epsilon$, $Y$, and $\textrm{Emb}()$. These elements describe the original data, the underlying signal, and the embedding of the signal in data space, respectively. When simulating examples, it’s natural to start with the underlying signal $Y$ then construct $Z+\epsilon$ by attaching extra dimensions and adding Gaussian noise. The $\textrm{Emb}()$ function is then given by
$\textrm{Emb}(y)=(y,0,\ldots,0),$
so that
@@
Practical examples are more tricky because we do not have the luxury of first defining $Y$. Instead, we are given the data $Z+\epsilon$ from which we must extract $Y$, or at least our best estimate. This process is left to the researcher and should be based on a priori knowledge of the data. If there is no specific signal of interest, a more general approach can be taken. We used a PCA projection of the data to represent the signal,
$Y=\textrm{PCA}_{r}(Z+\epsilon),$
where $r$ is the dimension of the projection. For a reasonable number of dimensions $r$, we would expect the first $r$ principal components to contain most of the signal, while excluding most of the noise. Another advantage to using PCA is it gives rise to a natural $\textrm{Emb}()$ function — the PCA inverse transform. If $Y$ is centered, then we may define
@@
Figure 3a: Shepard Goodness vs. Trustworthiness (Links)
@@ 4.1 Simulated Examples
We first looked at simulated examples with explicitly defined signal structures – three low-dimensional examples (Figure 2) and one high-dimensional example.
@@
For the three low-dimensional examples, the signal $Y$ consisted of 500 points in three dimensions. The dataset $Z+\epsilon$ was constructed from $Y$ by adding seven superfluous dimensions and isotropic Gaussian noise to all 10 dimensions. We then ran t-SNE using the $R$ package Rtsne with perplexities ranging from 10 to 160. For each value of perplexity, we ran the algorithm 20 times to mimic the ordinary t-SNE workflow. If we were to disregard the distinction between signal and noise, a plot of $f_{\textrm{Shep}}(Z+\epsilon,X)$ vs. $f_{\textrm{trust}}(Z+\epsilon,X)$ could be used to calibrate perplexity. To avoid overfitting the noise, a plot of $f_{\textrm{Shep}}(Y,X)$ vs. $f_{\textrm{trust}}(Y,X)$ should be used instead. See Figure 3a for plots of the links example. Both plots depict an increase in global performance as perplexity increases. Both plots also depict an increase in local performance followed by a decrease as perplexity increases. This trend is more evident when we plot trustworthiness vs. perplexity (Figure 3b).
Figure 3b: Trustworthiness vs. Perplexity (Links)
Figure 4: Trustworthiness vs. Perplexity (High-Dimensional Clusters)
Past a certain perplexity, local performance begins to deteriorate with increasing perplexity, exhibiting the trade-off between global and local performance. This cutoff point, however, varies between the two plots. When comparing against the original data, a perplexity of 30 maximizes trustworthiness, which is consistent with the original authors’ suggestion of 5 to 50 for perplexity [17]. When comparing against the signal, a perplexity of 110 maximizes trustworthiness. We hypothesize t-SNE tends to overfit the noise when the perplexity is too low. Intuitively, small perplexities are more affected by slight perturbations of the data when only considering small neighborhoods around each point, leading to unstable representations. Conversely, larger perplexities lead to more stable representations that are less affected by noise. The other two low-dimensional simulated examples exhibit a similar trend (See Table 1 and Supplementary Information).
Figure 5: Scree Plot for CyTOF Dataset
Figure 6a: Shepard Goodness vs. Trustworthiness (CyTOF)
Figure 6b: Trustworthiness vs. Perplexity (CyTOF)
The high-dimensional example was constructed by generating seven 10-dimensional Gaussian clusters with varying centers. More specifically, the signal $Y$ contained 210 points in 10 dimensions. The dataset $Z+\epsilon$ was constructed from $Y$ by adding 50 superfluous dimensions and isotropic Gaussian noise to all 60 dimensions. Again, local performance peaked at different perplexities when changing the frame of reference (Figure 4). When comparing against the original data, trustworthiness was maximized at a perplexity of 40. When comparing against the underlying signal, trustworthiness was maximized at a perplexity of 50.

### v5

@@ Abstract
The goal of dimension reduction tools is to construct a low-dimensional representation of high-dimensional data. These tools are employed for a variety of reasons such as noise reduction, visualization, and to lower computational costs. However, there is a fundamental issue that is discussed in other modeling problems that is often overlooked in dimension reduction — overfitting. In the context of other modeling problems, techniques such as feature-selection, cross-validation, and regularization are employed to combat overfitting, but rarely are such precautions taken when applying dimension reduction. Prior applications of the two most popular non-linear dimension reduction methods, t-SNE and UMAP, fail to acknowledge data as a combination of signal and noise when assessing performance. These methods are typically calibrated to capture the entirety of the data, not just the signal. In this paper, we demonstrate the importance of acknowledging noise when calibrating hyperparameters and present a framework that enables users to do so. We use this framework to explore the role hyperparameter calibration plays in overfitting the data when applying t-SNE and UMAP. More specifically, we show previously recommended values for perplexity and n_neighbors are too small and overfit the noise. We also provide a workflow others may use to calibrate hyperparameters in the presence of noise.
@@ 1 Introduction
In recent years, non-linear dimension reduction techniques have been growing in popularity due to their usefulness when analyzing high-dimensional data. Biologists use these techniques for a variety of visualization and analytic purposes, including exposing cell subtypes [1], checking for batch effects [2], and visualizing the trajectories of differentiating cells [3]. The most popular non-linear dimension reduction methods are t-distributed Stochastic Neighbor Embedding (t-SNE, [4]) and Uniform Manifold Approximation and Projection (UMAP, [5]). Both methods have been applied to various types of data within biology ([1], [6], [7]).
@@
Since the introduction of t-SNE and UMAP, hyperparameter calibration has proven to be a difficult task. The most crucial hyperparameters, t-SNE’s perplexity and UMAP’s n_neighbors, control how large a neighborhood to consider around each point when determining its location in low dimension. Calibration is so troublesome, that perplexity-free versions of t-SNE have been proposed [8]. It is also an extremely important task, since both methods are known to produce unfaithful results when mishandled [9]. For t-SNE, the original authors suggested perplexities between 5 and 50 [4], while recent works have suggested perplexities as large as one percent of the sample size [10]. [11] studied the inverse relationship between perplexity and Kullback-Leibler divergence to design an automatic calibration process that “generally agrees with experts’ consensus.” For UMAP, the original authors make no recommendation for optimal values of n_neighbors, but their implementation defaults to n_neighbors = 15 [5]. Manual tuning of perplexity and n_neighbors requires a deep understanding of the t-SNE and UMAP algorithms, as well as a general knowledge of the data’s structure.
The primary purpose of dimension reduction is to simplify data in a way that eliminates superfluous or nonessential information, i.e. noise. Each dimension reduction method does this slightly differently, but most require hyperparameter calibration. For example, the classical linear method, PCA, requires tuning of the number of principal components. A more contemporary method in biology, PHATE (Potential of Heat-diffusion for Affinity-based Trajectory Embedding) [3], requires tuning of a hyperparameter named diffusion time scale $t$. PHATE represents the structure of the data by computing local similarities then walking through the data using a Markovian random-walk diffusion process. $t$ determines the number of steps taken in a random walk and “provides a tradeoff between encoding local and global information in the embedding” [3]. Perplexity and n_neighbors serve the same purpose in their respective algorithms. Hence, we believe t-SNE and UMAP are capable of handling noise, but naïve calibrations that disregard noise often result in overfitting.
To assess dimension reduction performance in the presence of noise, we must acknowledge noise during the evaluation process. When the data’s structure is available, we can visualize the results and choose the representation that best captures the hypothesized structure. In supervised problems, for example, we look for low-dimensional representations that cluster according to the class labels. For unsupervised problems, however, the structure is often unknown, so we cannot visually assess each representation. In these cases, we must resort to quantitative measures of performance to understand how well the low-dimensional representation reproduces the high-dimensional data. While this strategy is heavily discussed in the machine learning literature, many prior works disregard the possibility of overfitting when quantitatively measuring performance.
@@
In this paper, we present a framework for studying dimension reduction methods in the presence of noise (Section 3). We then use this framework to calibrate t-SNE and UMAP hyperparameters in both simulated and practical examples to illustrate how the disregard of noise leads to miscalibration (Section 4). We also discuss how other researchers may use this framework in their own work (Section 5) and present a case study that walks the reader through the application of the framework to a modern data set (Section 6).
@@ 2 Background
Thus for each point $x_{i}$, we have a probability distribution $P_{i}$ that quantifies the similarity of $x_{i}$ to every other point. The scale of the Gaussian kernel, $\sigma_{i}$, is chosen so that the perplexity of the probability distribution $P_{i}$, in the information theory sense, is equal to a pre-specified value also named perplexity,
@@
Intuitively, perplexity controls how large a neighborhood to consider around each point when approximating the topological structure of the data. As such, it implicitly balances attention to local and global aspects of the data with high values of perplexity placing more emphasis on global aspects. For the sake of computational convenience, t-SNE assumes the directional similarities are symmetric by defining
@@
The primary downsides of t-SNE are its inherent randomness, unintuitive results, and sensitivity to hyperparameter calibration. The minimization of KL divergence is done using gradient descent methods with incorporated randomness to avoid stagnating at local minima. As a result, the output differs between runs of the algorithm. Hence, the traditional t-SNE workflow often includes running the algorithm multiple times at various perplexities before choosing the best representation. t-SNE is also known to produce results that are not faithful to the true structure of the data, even when calibrated correctly. For example, cluster sizes and inter-cluster distances aren’t always consistent with the original data [13]. Such artifacts of the t-SNE algorithm can be confused for significant structures by inexperienced users.
@@ 2.2 UMAP
Uniform Manifold Approximation and Projection (UMAP, [5]) is another nonlinear dimension reduction method that has been rising in popularity. Originally introduced as a more computationally efficient alternative to t-SNE, UMAP is a powerful tool for visualizing high-dimensional data that requires user calibration. While its underlying ideology is completely different from that of t-SNE, the UMAP algorithm is very similar architecturally to the t-SNE algorithm — high-dimensional similarities are computed and the resulting representation is the set of low-dimensional points whose low-dimensional similarities best match the high-dimensional similarities. See [5] for details. The largest difference is UMAP’s default initialization process. UMAP uses Laplacian eigenmaps to initialize the low-dimensional representation, which is then adjusted to minimize the cost function. Most t-SNE implementations use PCA during the initialization process. The initialization process is the primary benefit of the default implementation of UMAP, but t-SNE and UMAP have been shown to perform similarly with identical initializations [7]. Modern implementations of both algorithms are also comparable in speed.
UMAP shares similar disadvantages with t-SNE. It can create unfaithful representations that require experience to interpret and is sensitive to hyperparameter calibration [14].
@@ 3.1 Dimension Reduction Framework
Prior works quantitatively measure how well low-dimensional representations match the high-dimensional data. However, if we consider data as a composition of signal and noise, we must not reward capturing the noise. Therefore, we should be comparing the low-dimensional representation against the signal underlying our data, rather than the entirety of the data.
Suppose the underlying signal of our data is described by an $r$-dimensional matrix $Y\in\mathbb{R}^{n\times r}$. In the context of dimension reduction, the signal is often lower dimension than the original data. Let $p\geq r$ be the dimension of the original data set, and let $\textrm{Emb}:\mathbb{R}^{r}\to\mathbb{R}^{p}$ be the function that embeds the signal in data space. Define $Z=\textrm{Emb}(Y)$ to be the signal embedded in data space. We then assume the presence of random error. The original data can then be modeled by $Z+\epsilon\textrm{ for }\epsilon\sim N_{p}(0,\Sigma)$. The dimension reduction method $\varphi$ is applied to $Z+\epsilon$ to get a low-dimensional representation $X\in\mathbb{R}^{n\times q}$. See Figure 1.
${Z+\epsilon}$${X}$${Y}$${Z}$$\scriptstyle{\varphi}$EmbDimension (low to high)
@@ 3.2 Reconstruction Error Functions
The remaining piece is a procedure for measuring dimension reduction performance. Suppose we have a reconstruction error function $f(D_{1},D_{2})$ that quantifies how well the data set $D_{2}$ represents the data set $D_{1}$. Prior works like [9], [10] , [15], and [16] use various reconstruction error functions to quantify performance; only, they study $f(Z+\epsilon,X)$ to measure how well the constructed representation $X$ represents the original data $Z+\epsilon$. We argue it is more appropriate to compare $X$ against the signal $Y$ by examining $f(Y,X)$.
Prior works in dimension reduction have suggested various quantitative metrics for measuring dimension reduction performance. In line with recent discussions of perplexity ([10] and [15]), we employ two different metrics — one that measures local performance and one that measures global performance.
@@
Local performance is the primary concern when applying t-SNE and UMAP, so our experiments focus on maximizing trustworthiness.
For global performance, we use Shepard goodness [16]. Shepard goodness is the Spearman correlation, a rank-based correlation, between high and low-dimensional inter-point distances,
$f_{\textrm{Shep}}(D_{1},D_{2})=\sigma_{\textrm{Spearman}}(||z_{i}-z_{j}||,||\varphi(z_{i})-\varphi(z_{j})||).$
@@ 3.3 Using this framework
When using this framework to model examples, three components must be specified: $Z+\epsilon$, $Y$, and $\textrm{Emb}()$. These elements describe the original data, the underlying signal, and the embedding of the signal in data space, respectively. When simulating examples, it’s natural to start with the underlying signal $Y$ then construct $Z+\epsilon$ by attaching extra dimensions and adding Gaussian noise. The $\textrm{Emb}()$ function is then given by $\textrm{Emb}(y)=(y,0,\ldots,0)$ so that
@@
Practical examples are more tricky because we do not have the luxury of first defining $Y$. Instead, we are given the data $Z+\epsilon$ from which we must extract $Y$, or at least our best estimate. This process is dependent on the specifics of the problem and should be based on a priori knowledge of the data. If there is no specific signal of interest, a more general approach can be taken. We used a PCA projection of the data to represent the signal, $Y=\textrm{PCA}_{r}(Z+\epsilon)$, where $r$ is the dimension of the projection. For a reasonably chosen $r$, we expect the first $r$ principal components to contain most of the signal, while excluding most of the noise. Another advantage to using PCA is it gives rise to a natural $\textrm{Emb}()$ function — the PCA inverse transform. If $Y$ is centered, then we may define
@@
Figure 3: Trustworthiness vs. Perplexity (Links $sd=1$)
Figure 4: Trustworthiness-Maximizing Representations (Links $sd=1$)
@@ 4.1 Simulated Examples
We first looked at simulated examples with explicitly defined signal structures – three low-dimensional examples (Figure 2) and one high-dimensional example. The links example and the high-dimensional example are explored here. See Table 1 and the Supporting Information (SI.1 and SI.2) for the other simulated examples.
4.1.1 Links Data Set
For the links example, the signal $Y$ consisted of two interlocked circles, each containing 250 points, embedded in three dimensions. $Z+\epsilon$ was constructed by adding seven superfluous dimensions and isotropic Gaussian noise. Various degrees of noise were tested ($sd=0.5,1,1.5,2,2.5,3$).
@@ 4.1.1 Links Data Set
t-SNE was run using the $R$ package Rtsne [19] at varying perplexities. For each perplexity, the algorithm was run 40 times to mimic the ordinary t-SNE workflow. If the distinction between signal and noise were disregarded, a plot of $f_{\textrm{trust}}(Z+\epsilon,X)$ vs. perplexity could be used to maximize local performance. To avoid overfitting the noise, a plot of $f_{\textrm{trust}}(Y,X)$ vs. perplexity should be used instead. See Figure 3 for examples of these plots for the $sd=1$ case. Both plots depict an increase in local performance followed by a decrease as perplexity increases. This cutoff point, however, varies between the two plots. When comparing against the original data, the trustworthiness-maximizing representation was constructed with a perplexity of 40, which is consistent with the original authors’ suggestion of 5 to 50 for perplexity [4]. When comparing against the signal, the trustworthiness-maximizing representation was constructed with a perplexity of 80.
With the signal structure known, we are also able to visually assess the trustworthiness-maximizing representations. Figure 4 shows the trustworthiness-maximizing representations for the $sd=1$ case. Notice the larger perplexity was able to successfully separate the circles in the presence of noise, while the smaller perplexity was not. By using the signal as the frame of reference, our framework correctly rewarded the representation that was able to successfully separate the two links.
The same pattern held true for other levels of noise. The optimal perplexity was consistently larger when comparing against the signal, rather than the original data (Figure 5).
Figure 5: Optimal Perplexity (Links)
These results suggest larger perplexities perform better in the presence of noise, both quantitatively and qualitatively. We hypothesize t-SNE tends to overfit the noise when the perplexity is too small. Intuitively, small perplexities are more affected by slight perturbations of the data when only considering small neighborhoods around each point, leading to unstable representations. Conversely, larger perplexities lead to more stable representations that are more robust to noise.
Figure 6: Trustworthiness vs. Perplexity (High-Dimensional Clusters $sd=3$)
Figure 7: Trustworthiness-Maximizing Representations (High-Dimensional Clusters $sd=3$)
Figure 8: Optimal Perplexity (High-Dimensional Clusters)
Figure 9: Scree Plot for scRNA-seq Data Set
Figure 10: Trustworthiness vs. Perplexity for r = 5 (scRNA-seq)
Figure 11: Trustworthiness vs. Perplexity for r = 5 (scRNA-seq)
4.1.2 High-Dimensional Clusters
The signal $Y$ consisted of seven Gaussian clusters, each containing 50 points, in seven dimensions. The clusters were drawn from multivariate normal distributions with mean $10e_{i}$ and random diagonal covariance matrices, where $e_{i}$ is the $i^{\textrm{th}}$ standard basis vector. The data set $Z+\epsilon$ was constructed from $Y$ by adding 53 superfluous dimensions and isotropic Gaussian noise to all 60 dimensions. Various degrees of noise were tested $(sd=2,2.5,3,3.5,4,4.5)$.
When $sd=3$, local performance peaked at different perplexities when changing the frame of reference (Figure 6). When comparing against the original data, trustworthiness was maximized at a perplexity of 55. When comparing against the signal, trustworthiness was maximized at a perplexity of 60. See Figure 7 for the trustworthiness-maximizing representations. Visually, both representations maintain the original clustering to some extent, but the higher-perplexity representation shows less mixing between the clusters and had a larger average silhouette width (0.178) than the lower-perplexity representation (0.121). This suggests the higher-perplexity representation better maintained the original clustering.
Figure 8 shows the optimal perplexities for different levels of noise. Again, the trustworthiness-maximizing perplexity was larger when comparing against the signal for all levels of noise.